OpenStudy (anonymous):

Solve for x: 3/4 x + 5/8 = 4x

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

5/8 = (4-3/4)x 5/8 / 13/4 = x x = 5/26 x = 0.192308 Proved with http://www.hackmath.net/en/calculator/solving-system-of-linear-equations?input=+3%2F4+x+%2B+5%2F8+%3D+4x&submit=Calculate